OpenStudy (anonymous):

Help me solve (^6 sqrt of x)^3?

OpenStudy (anonymous):

It originally equals x^1/3

OpenStudy (anonymous):

But the question is: \[(\sqrt[6]{x})^3\]

OpenStudy (anonymous):

It's the parenthesis that throws me off..

terenzreignz (terenzreignz):

You can, of course, express everything as exponents first. That makes easier (to figure out) In particular, sixth root simply means an exponent of 1/6. \[\LARGE \sqrt[6]x= \color{blue}{x^{\frac16}}\] In effect, \[\LARGE (\sqrt[6]x)^3 = \left(\color{blue}{x^{\frac16}}\right)^3\] Which may now be simplified according to the laws of exponents...

hero (hero):

I already wrote out the rule to you @whalexnuker

hero (hero):

It's the same format

OpenStudy (anonymous):

Perhaps that might help me better... I didn't think of simplifying it from that. Lol <_> I'm still getting a result of sqrt of x. the parenthesis are still throwing me off. Solving it on my own (because i was using a program), i would like to say you'd multiply everything on the inside with the 3? Yeah, hero. Let me look at that because the problem is the same LOL. i just noticed. But since I have no C, it's just..... x^3/6?

OpenStudy (anonymous):

Wait, sorry. I don't have A. but the equation is the same

OpenStudy (anonymous):

it just the parenthesis that still throws me off... I realize its the same as the other but I don't have A so nothing is being multiplied..

OpenStudy (anonymous):

is it x^1/2? am I dividing the 3 with the 6?

OpenStudy (anonymous):

riiiggghhhtt? >.> @Hero

hero (hero):

correct, but @terenzreignz helped you with this. Give the medal to him.

OpenStudy (anonymous):

Done :D I couldn't have done it without that simplified form. I would have been stuck on this forever. Thanks.

OpenStudy (anonymous):

What about ^5 sqrt (x) * (x^2/5) * (x^3/5) * (x^4/5) @terenzreignz >.> since the denominators are the same outside of the square root, i can add the exponents together, yes? 2+3+5 = x^10/5 ^5sqrt(x) x^10/5

OpenStudy (anonymous):

With the 5 that's on the input, does that get divided by the 10/5 or something?

OpenStudy (anonymous):

it equals x^2 but I'm trying to figure out how.

OpenStudy (anonymous):

@bhaskar.pathk Are you just going to sit on my question....? you've been here for a bit...

terenzreignz (terenzreignz):

Sorry, was spaced out... hang on...

terenzreignz (terenzreignz):

But didn't you already figure it out? Adding the exponents gives 1/5 + 2/5 + 3/5 + 4/5 = 10/5 = 2

terenzreignz (terenzreignz):

Not forgetting, of course, that \[\LARGE \sqrt[5]x = x^{\frac15 }\]

OpenStudy (anonymous):

@whalexnuker : sorry was facing some network prblm

OpenStudy (anonymous):

I think I solved half of it :p I got the 10/5 which equals 2 but I still have the ^5 sqrt of (x). So, right now, it's ^5sqrt of (x) 10/5 (or 2). I dunno what to do with that "^5" since the square root is only covering the X and not the 2

terenzreignz (terenzreignz):

like this: \[\Large \sqrt[5]x \cdot x^{\frac25}\cdot x^{\frac35}\cdot x^{\frac45}\]

terenzreignz (terenzreignz):

But because of this fact: \[\LARGE \sqrt[5]x = x^{\frac15 }\] That expression on top is just equal to \[\Large x^{\frac15} \cdot x^{\frac25}\cdot x^{\frac35}\cdot x^{\frac45}\] And now, the exponents may be easily added to get the desired result.

terenzreignz (terenzreignz):

like this: \[\Large \sqrt[5]x \cdot x^{\frac25}\cdot x^{\frac35}\cdot x^{\frac45}\]

OpenStudy (anonymous):

!! oh! See, I forgot to switch it again XD I can now just add them! Awesome! and THAT gives me x^10/5 = x^2 I'd give you medals if I could. In fact, let me scroll through your profile and I will :D

terenzreignz (terenzreignz):

No, there's no need for that, as long as you got it, that's enough :)

OpenStudy (anonymous):

It's just duh moments >.> My whole lesson is about converting them like that and I let it completely slip my mind lmao. Thanks so much.

terenzreignz (terenzreignz):

When in doubt, turn everything into exponents, fractional or not, they're generally easier to work with.

terenzreignz (terenzreignz):

...than that nasty radical sign \(\Large \sqrt{\qquad}\)

terenzreignz (terenzreignz):

I need to sign off now... good luck with the rest (if any) ^_^ ---------------------- Terence out